# Stability analysis through folds: An end-loaded elastica with a lever arm

### v3

@@ 2.2 Varying Parameter at the Fixed-end
In this section, we analyze the behavior of $-\langle\mathcal{S}\ddot{\boldsymbol{\zeta}},\dot{\boldsymbol{\zeta}}\rangle$ and consequently $\frac{d}{d\tau}\delta^{2}J$ when $\xi_{o}$ is varied. At the fold, we have $\dot{\xi}_{0}=0$, $\dot{\boldsymbol{\zeta}}(0)=0$ and $\left(\mathbf{P}\dot{\boldsymbol{\zeta}}^{\prime}+\mathbf{C}^{T}\dot{\boldsymbol{\zeta}}+\mathbf{B}\dot{\boldsymbol{\zeta}}\right)_{s=l}=0$. Using the relations (20) and recalling that $\mathbf{P}=\mathbf{P}^{T}$, we deduce
$-\langle\mathcal{S}\ddot{\boldsymbol{\zeta}},\dot{\boldsymbol{\zeta}}\rangle$ $=\left(\mathbf{P}\dot{\boldsymbol{\zeta}}^{\prime}+\mathbf{C}^{T}\dot{\boldsymbol{\zeta}}+\mathbf{B}\dot{\boldsymbol{\zeta}}\right)\cdot\ddot{\boldsymbol{\zeta}}\bigg{|}_{s=0}+\left(\mathbf{P}\ddot{\boldsymbol{\zeta}}^{\prime}+\mathbf{C}^{T}\ddot{\boldsymbol{\zeta}}+\mathbf{B}\ddot{\boldsymbol{\zeta}}\right)\cdot\dot{\boldsymbol{\zeta}}\bigg{|}_{s=l},$
$=\ddot{\xi}_{0}\left(\frac{\partial\boldsymbol{\zeta}_{o}}{\partial\xi_{0}}\cdot\mathbf{P}\dot{\boldsymbol{\zeta}}^{\prime}\right)\Bigg{|}_{s=0}+\left(\mathbf{C}^{T}\dot{\boldsymbol{\zeta}}+\mathbf{B}\dot{\boldsymbol{\zeta}}\right)\cdot\ddot{\boldsymbol{\zeta}}\bigg{|}_{s=0}+\left(\mathbf{P}\ddot{\boldsymbol{\zeta}}^{\prime}+\mathbf{C}^{T}\ddot{\boldsymbol{\zeta}}+\mathbf{B}\ddot{\boldsymbol{\zeta}}\right)\cdot\dot{\boldsymbol{\zeta}}\bigg{|}_{s=l},$
@@
For the function $\mathbf{P}=\mathcal{L}_{\boldsymbol{\zeta}^{\prime}\boldsymbol{\zeta}^{\prime}}$, which is a symmetric matrix, we can write
@@ 2.3 Varying Parameter at the Free-end
Now consider the other case, where the parameter is at the free end $s=l$, namely $\xi_{l}$. Let us go back to the expression for $-\langle\mathcal{S}\ddot{\boldsymbol{\zeta}},\dot{\boldsymbol{\zeta}}\rangle$ given by (22). At the fold $\dot{\xi}_{l}=0$, $\dot{\boldsymbol{\zeta}}(0)=0$ and $\left(\mathbf{P}\dot{\boldsymbol{\zeta}}^{\prime}+\mathbf{C}^{T}\dot{\boldsymbol{\zeta}}+\mathbf{B}\dot{\boldsymbol{\zeta}}\right)_{s=l}=\mathbf{0}$, and hence we have
@@
Additionally, using the boundary conditions given by (21b) results in
@@ 3 Planar Elastica
To test the developed theory in  section 2, we consider a special Euler elastica and analyze stability of the associated equilibria. An inextensible and unshearable planar massless uniform elastic rod with a bending stiffness $K$, which is straight in its undeformed configuration, is clamped at one end and connected to a dead load through a massless rigid lever arm at the other, as shown in Figure 2. Let $\{\mathbf{e_{x}},\mathbf{e_{y}},\mathbf{e_{z}}\}$ denote a fixed right-handed orthonormal laboratory frame. The ends of the rod are defined by the coordinates $s=0$ and $s=l$ in its fixed undeformed configuration. The lever arm is rigidly fixed to the tip $s=l$ so that it forms an angle $\psi$ with the tangent of the tip $s=l$. The dead load exerts a force $\mathbf{F}$ in the $\mathbf{e_{x}}-\mathbf{e_{y}}$ plane and a moment $\Psi\times\mathbf{F}$ due to an offset of the loading at the tip $s=l$. Here, $\Psi$ is the lever arm of the load in the fixed frame. Let $\mathbf{d}_{r}$ and $\mathbf{d}_{t}$ denote the local directors along the radial and tangential directions, respectively, of the centerline at the tip $s=l$, given by
$\mathbf{d}_{t}=-\sin\theta(l)\mathbf{e}_{x}+\cos\theta(l)\mathbf{e}_{y},\qquad\mathbf{d}_{r}=\cos\theta(l)\mathbf{e}_{x}+\sin\theta(l)\mathbf{e}_{y}$ (31)
Then, $\Psi$ is given by $\Delta_{r}\mathbf{d_{r}}+\Delta_{t}\mathbf{d_{t}}$, where the constants $\Delta_{r}$ and $\Delta_{t}$ are the respective components. If $\Delta$ denotes the length of the arm and makes an angle $\psi$ with the tangent at the tip $s=l$, then $\Delta_{t}=\Delta\cos\psi$ and $\Delta_{r}=\Delta\sin\psi$. Let $\theta(s)$ be be the angle between the tangent of the rod centerline and the vertical. The direction of the load $\mathbf{F}$ with respect to vertical laboratory frame is denoted by $\alpha$. The position vector of the rod centerline $\mathbf{r}(s)$ in the deformed configuration is represented using
$\mathbf{r}(s)=x(s)\mathbf{e_{x}}+y(s)\mathbf{e_{y}},$ (32)
where $x(s)$ and $y(s)$ are the components in the fixed frame. The inextensibility and unshearability condition restricts the tangent of the rod in the deformed configuration to satisfy:
@@
Using the principles of differential geometry, the curvature of the rod’s centerline can be easily expressed as $\theta^{\prime}$. Let $\mathbf{m}(s)$ denote the internal moment vector along the rod. We consider a simple linear constitutive law where the planar bending moment in the rod is related to the curvature $\theta^{\prime}$ through
@@
The clamped end is constrained to rotate about $\mathbf{e}_{z}$- axis. Since it is a planar case, the moment vector $\mathbf{m}$ is always oriented perpendicular to the plane $(\mathbf{m}\cdot\mathbf{e}_{x}=\mathbf{m}\cdot\mathbf{e}_{y}=0)$. The total energy stored in the system is the sum of the bending energy and the work done by the external force, satisfying the inextensibility and unshearability constraint (33)
$\mathcal{L}=\int_{0}^{l}\frac{K}{2}\theta^{\prime 2}(s)+n_{x}(x^{\prime}+\sin\theta)+n_{y}(y^{\prime}-\cos\theta)ds-\mathbf{F}\cdot(\mathbf{r}(l)+\Psi(l)),$ (36)
where $n_{x}$ and $n_{y}$ are the corresponding Lagrange multipliers and denote the internal forces along $\mathbf{e}_{x}$ and $\mathbf{e}_{y}$ axes, respectively. We use the vector $\mathbf{n}=[n_{x},n_{y}]^{T}$ to represent them. The equilibrium equations are obtained by applying the Euler-Lagrange equations on (36)
@@
$K\theta^{\prime}(l)-\left(\Psi\times\mathbf{F}\right)\cdot\mathbf{e}_{z}=0,$ (38b)
where $F_{x}$ and $F_{y}$ are the components of $\mathbf{F}$ along $\mathbf{e}_{x}$ and $\mathbf{e}_{y}$ directions, respectively. Indeed, the Euler-Lagrange equations given by (37) are equivalent to planar versions of force and moment balance in elastic rods [14]
@@
We also represent the applied load $\mathbf{F}$ using its magnitude and direction $\alpha$ as
@@
In this apparatus, the parameters that describe the Dirichlet boundary conditions, such as $\theta_{o}$ and those that characterize the natural boundary conditions such as $|\mathbf{F}|$, $\Delta$, $\alpha$ and $\psi$ can be varied. We perform several numerical experiments using parameter continuation [1] and employ the proposed distinguished bifurcation diagrams to determine the stability properties of the generated parameter-dependent equilibria.
@@ 3.1 Bifurcation Analysis
The distinguished bifurcation diagram gives information on the changes in stability. So, the stability of atleast one equilibrium along a branch must be established before proceeding. Hence, bifurcation analysis is performed on the current apparatus to evaluate the indices. The buckling characteristics of the elastica are well known; now, let us examine how the lever arm influences this behavior. For this analysis, we set $\alpha=0$ and $\psi=0$, and consequently we obtain $\lambda_{x}=0$ and $\lambda_{y}=-\|\mathbf{F}\|$. We obtain the classical elastica when $\Delta=0$. Before proceeding further, we non-dimensionalize the problem (37b) by substituting $\bar{s}=\frac{s}{l}$, $\epsilon=\frac{\Delta}{l}$ and $P=\frac{\|\mathbf{F}\|l^{2}}{K}$ that fetches
@@
$\theta(0)$ $=0,\qquad\theta^{\prime}(1)-P\epsilon\sin\theta=0$ (41b)
@@
$\sqrt{P}\cos\sqrt{P}-P\epsilon\sin\sqrt{P}=0,\implies\cot\sqrt{P}-\sqrt{P}\epsilon=0,\qquad\sqrt{P}\neq n\pi.$ (44)
Figure 3: Plot illustrating the solutions of $\cot\sqrt{P}-\sqrt{P}\epsilon=0$ for different values of $\epsilon$, along with the local bifurcation characteristics associated with $\epsilon=+/-0.25$ shown at top.
The nontrivial solutions satisfying this relation that correspond to bifurcations are depicted in Figure 3. The critical buckling modes of elastica are the solutions pertaining to $\epsilon=0$ are $P^{o}_{cr}=\frac{(2k-1)^{2}\pi^{2}}{4},k=1,\dots,\infty$. Let $P_{cr}^{+}$ represent the critical load values pertaining to positive $\epsilon$, while $P_{cr}^{-}$ represent pertaining to negative $\epsilon$. Note that for each $k$ th mode, $(k-1)^{2}\pi^{2}<P_{cr,k}^{+}<\left(\frac{(2k-1)}{2}\right)^{2}\pi^{2}$ and $\left(\frac{(2k-1)}{2}\right)^{2}\pi^{2}<P_{cr,k}^{-}<k^{2}\pi^{2}$, where $k=1,2,\dots,\infty$. It can be easily shown that these points correspond to supercritical pitchfork bifurcations, after a similar analysis in  [14, pp, 174-177]. The linearized equation can be rewritten as an eigenvalue problem:
@@
$\delta\theta^{\prime\prime}-\left(P+\mu\right)\delta\theta=0,\qquad\delta\theta(0)=0,\qquad\delta\theta^{\prime}(1)-P\epsilon\delta\theta(1)=0.$ (46)
Typically, at a supercritical bifurcation [4], the number of negative eigenvalues increases by one with each successive bifurcation. In our setup, the index of the equilibria for load $P_{cr,k}^{+/-}$ is $k$. For a very high $\epsilon$ value, i.e., when the $\sqrt{P}\epsilon$ line approaches the vertical line, the value of $P_{cr}^{+}$ approaches zero, suggesting that bifurcations occur at a very small load when the arm is long. On the other hand, $P_{cr}^{-}$ approaches $\pi^{2}$ for higher loads, which indicates the critical loads for the case of fixed-fixed ends. The trivial solutions are stable only for loads below $P_{cr}^{+/-}$. Consider two instances of loads which correspond to first mode and second mode, say $\pi^{2}/4$ and $49\pi^{2}/$ 4. These chosen straight equilibria correspond to index one and two respectively, while the negative $\epsilon$ leads to index zero and index one respectively. We use these equilibria as initial solutions for numerical continuation in the subsequent sections.
@@ 4 Numerical Examples
The distinguished bifurcation diagrams facilitate the determination of the index change at a fold in the varying parameter. If there are no folds, the stability remains the same for all solutions along the branch. At folds, the stability transitions usually happen, and they were evaluated by examining the manner in which the fold occurs as depicted in Figure 3. We use the straight trivial solutions from previous section as initial solutions to generate the family of equilibria by varying parameters. Subsequently, we assess the stability of the resulting equilibria. Numerical continuation is performed using AUTO-07p [39], which employs pseudo-arclength continuation [1], and is capable of detecting singularities like folds and bifurcations as the parameter is varied. In these examples, the analysis is restricted solely to simple folds.

### v5

@@ 2.2 Varying Parameter at the Fixed-end
In this section, we analyze the behavior of $\langle\mathcal{S}\ddot{\boldsymbol{\zeta}},\dot{\boldsymbol{\zeta}}\rangle$ and consequently $\frac{d}{d\tau}\delta^{2}J$ when $\xi_{0}$ is varied. At the fold $\dot{\xi}_{0}=0$, (14) leads to $\dot{\boldsymbol{\zeta}}(0)=\mathbf{0}$ and $\left(\mathbf{P}\dot{\boldsymbol{\zeta}}^{\prime}+\mathbf{C}^{T}\dot{\boldsymbol{\zeta}}+\mathbf{B}\dot{\boldsymbol{\zeta}}\right)_{s=l}=\mathbf{0}$. Then, (22) simplifies to
$-\langle\mathcal{S}\ddot{\boldsymbol{\zeta}},\dot{\boldsymbol{\zeta}}\rangle$ $=\left(\mathbf{P}\dot{\boldsymbol{\zeta}}^{\prime}+\mathbf{C}^{T}\dot{\boldsymbol{\zeta}}+\mathbf{B}\dot{\boldsymbol{\zeta}}\right)\cdot\ddot{\boldsymbol{\zeta}}\bigg{|}_{s=0}+\left(\mathbf{P}\ddot{\boldsymbol{\zeta}}^{\prime}+\mathbf{C}^{T}\ddot{\boldsymbol{\zeta}}+\mathbf{B}\ddot{\boldsymbol{\zeta}}\right)\cdot\dot{\boldsymbol{\zeta}}\bigg{|}_{s=l}.$
Using the relations (20) and recalling that $\mathbf{P}=\mathbf{P}^{T}$, we deduce
$-\langle\mathcal{S}\ddot{\boldsymbol{\zeta}},\dot{\boldsymbol{\zeta}}\rangle$ $=\ddot{\xi}_{0}\left(\frac{\partial\boldsymbol{\zeta}_{o}}{\partial\xi_{0}}\cdot\mathbf{P}\dot{\boldsymbol{\zeta}}^{\prime}\right)\Bigg{|}_{s=0}+\left(\mathbf{C}^{T}\dot{\boldsymbol{\zeta}}+\mathbf{B}\dot{\boldsymbol{\zeta}}\right)\cdot\ddot{\boldsymbol{\zeta}}\bigg{|}_{s=0}+\left(\mathbf{P}\ddot{\boldsymbol{\zeta}}^{\prime}+\mathbf{C}^{T}\ddot{\boldsymbol{\zeta}}+\mathbf{B}\ddot{\boldsymbol{\zeta}}\right)\cdot\dot{\boldsymbol{\zeta}}\bigg{|}_{s=l},$
@@
For the symmetric matrix $\mathbf{P}=\mathcal{L}_{\boldsymbol{\zeta}^{\prime}\boldsymbol{\zeta}^{\prime}}$, we can write
@@ 2.3 Varying Parameter at the Free-end
Now consider the other case, where the parameter is at the free end $s=l$, namely $\xi_{l}$. At the fold $\dot{\xi}_{l}=0$,  (14) leads to $\dot{\boldsymbol{\zeta}}(0)=0$ and $\left(\mathbf{P}\dot{\boldsymbol{\zeta}}^{\prime}+\mathbf{C}^{T}\dot{\boldsymbol{\zeta}}+\mathbf{B}\dot{\boldsymbol{\zeta}}\right)_{s=l}=\mathbf{0}$. As a result, the expression for $\langle\mathcal{S}\ddot{\boldsymbol{\zeta}},\dot{\boldsymbol{\zeta}}\rangle$ given by (22) turns to
@@
Furthermore, the boundary conditions specified in (21b) result in
@@ 3 Planar Elastica
To test the theory developed in section 2, we consider a special Euler elastica. In this section, we describe this apparatus and subsequently examine its straight equilibria and critical parameters.
3.1 Setup
@@
An inextensible and unshearable planar massless uniform elastic rod with a bending stiffness $K$, which is straight in its undeformed configuration, is clamped at one end and connected to a dead load through a massless rigid lever arm at the other end, as shown in Figure 2. Let $\{\mathbf{e}_{x},\mathbf{e}_{y},\mathbf{e}_{z}\}$ denote a fixed right-handed orthonormal laboratory frame with $\mathbf{e}_{z}$ pointing out of the plane. The ends of the rod are defined by the coordinates $s=0$ and $s=l$ in its fixed undeformed configuration. The lever arm is rigidly fixed to the tip $s=l$ so that it forms an angle $\psi$ with the tangent of the tip $s=l$. The dead load exerts a force $\mathbf{F}$ in the $\mathbf{e}_{x}-\mathbf{e}_{y}$ plane and a moment $\boldsymbol{\Psi}\times\mathbf{F}$ due to an offset of the load $\mathbf{F}$ at $s=l$. Here, $\boldsymbol{\Psi}$ is the lever arm of the load.
The position vector of the rod centerline $\mathbf{r}(s)$ in the deformed configuration is represented using
$\mathbf{r}(s)=x(s)\mathbf{e}_{x}+y(s)\mathbf{e}_{y}.$ (31)
Let $\theta(s)$ be be the angle between the tangent of the rod centerline and the vertical. Let $\mathbf{d}_{r}$ and $\mathbf{d}_{t}$ denote the local directors of the material cross-section and are represented by
$\mathbf{d}_{t}(s)=-\sin\theta(s)\mathbf{e}_{x}+\cos\theta(s)\mathbf{e}_{y},\qquad\mathbf{d}_{r}(s)=\cos\theta(s)\mathbf{e}_{x}+\sin\theta(s)\mathbf{e}_{y}.$ (32)
The inextensibility and unshearability of the rod constrains the director $\mathbf{d}_{t}$ along the tangent of the centerline resulting in
@@
Then, $\boldsymbol{\Psi}$ is given by $\Delta_{r}\mathbf{d}_{r}(l)+\Delta_{t}\mathbf{d}_{t}(l)$, where the constants $\Delta_{r}$ and $\Delta_{t}$ are the respective components. If $\Delta$ denotes the length of the arm and makes an angle $\psi$ with the tangent at the tip $s=l$, then $\Delta_{t}=\Delta\cos\psi$ and $\Delta_{r}=\Delta\sin\psi$. Using the principles of differential geometry, the curvature of the rod’s centerline can be easily expressed as $\theta^{\prime}$. Let $\mathbf{m}(s)$ denote the internal moment vector along the rod. We consider a simple linear constitutive law where the planar bending moment in the rod is related to the curvature $\theta^{\prime}$ through
@@
The clamped end is rotated about $\mathbf{e_{z}}$- axis. Since it is a planar case, the moment vector $\mathbf{m}$ is always oriented perpendicular to the plane $(\mathbf{m}\cdot\mathbf{e}_{x}=\mathbf{m}\cdot\mathbf{e}_{y}=0)$. The total energy stored in the system is the sum of the bending energy and the work done by the external force, satisfying the inextensibility and unshearability constraint (33)
$\mathcal{L}=\int_{0}^{l}\frac{K}{2}\theta^{\prime 2}+n_{x}(x^{\prime}+\sin\theta)+n_{y}(y^{\prime}-\cos\theta)ds-\mathbf{F}\cdot(\mathbf{r}(l)+\boldsymbol{\Psi}(l)),$ (36)
where $n_{x}$ and $n_{y}$ are the corresponding Lagrange multipliers and denote the internal forces along $\mathbf{e}_{x}$ and $\mathbf{e}_{y}$ axes, respectively. We represent them using the vector $\mathbf{n}=[n_{x},n_{y}]^{T}$. The equilibrium equations are obtained by applying the Euler-Lagrange equations (3) on (36)
@@
$K\theta^{\prime}(l)-\left(\boldsymbol{\Psi}\times\mathbf{F}\right)\cdot\mathbf{e}_{z}=0,$ (38b)
where $F_{x}$ and $F_{y}$ are the components of $\mathbf{F}$ along the $\mathbf{e}_{x}$ and $\mathbf{e}_{y}$ directions, respectively. In fact, the Euler-Lagrange equations given by (37) are equivalent to planar versions of the force and moment balance in elastic rods [14]
@@
The load vector $\mathbf{F}$ makes an angle $\alpha$ with respect to vertical as shown in Figure 2. The load $\mathbf{F}$ can be represented using its magnitude and direction $\alpha$ as
@@
In this apparatus, the parameters that describe the Dirichlet boundary conditions, such as $\theta_{o}$ and those that characterize the natural boundary conditions such as $|\mathbf{F}|$, $\Delta$, $\alpha$ and $\psi$ can be varied. We perform several numerical experiments using parameter continuation [1] and employ the proposed distinguished bifurcation diagrams to assess the stability of the equilibria.
@@ 3.2 Bifurcation Analysis
The distinguished bifurcation diagram provides information on changes in stability. So, the stability of atleast one equilibrium along a branch must be established before proceeding. For this reason, we conduct bifurcation analysis on the current apparatus and establish stability indices for some equilibria. The buckling characteristics of the elastica are well known; now, let us examine how the lever arm influences this behavior. For this analysis, we set $\alpha=0$ and $\psi=0$, and consequently we have $\lambda_{x}=0$ and $\lambda_{y}=-|\mathbf{F}|$. The classical elastica is obtained when $\Delta=0$. Before proceeding further, we non-dimensionalize (37b) by substituting $\bar{s}=\frac{s}{l}$, $\epsilon=\frac{\Delta}{l}$ and $P=\frac{\\ |\mathbf{F}|l^{2}}{K}$ that fetches
@@
$\theta(0)$ $=0,\qquad\theta^{\prime}(1)-P\epsilon\sin\theta=0.$ (41b)
@@
$\begin{split}&\sqrt{P}\cos\sqrt{P}-P\epsilon\sin\sqrt{P}=0,\\ &\implies\cot\sqrt{P}-\sqrt{P}\epsilon=0,\qquad\sqrt{P}\neq n\pi.\end{split}$ (44)
Figure 3: Plot illustrating the solutions of $\cot{\sqrt{P}}-\sqrt{P}\epsilon=0$ for different values of $\epsilon$, along with the local bifurcation characteristics associated with $\epsilon=+/-0.25$ shown at top. The family of configurations branching out at bifurcations are indicated.
The non-trivial solutions satisfying this relation that correspond to bifurcations are depicted in Figure 3. The critical buckling modes of elastica are the solutions pertaining to $\epsilon=0$ denoted as $P^{o}_{cr}=\frac{(2k-1)^{2}\pi^{2}}{4},k=1,2,\dots,\infty$. Let $P_{cr}^{+}$ represent the critical load values pertaining to positive $\epsilon$, while $P_{cr}^{-}$ represent pertaining to negative $\epsilon$. Note that for each $k$ th mode, $(k-1)^{2}\pi^{2}<P_{cr,k}^{+}<\left(\frac{2k-1}{2}\right)^{2}\pi^{2}$ and $\left(\frac{2k-1}{2}\right)^{2}\pi^{2}<P_{cr,k}^{-}<k^{2}\pi^{2}$, where $k=1,2,\dots,\infty$. It can be easily shown that these points correspond to supercritical pitchfork bifurcations, following a similar analysis in  [14, pp, 174-177]. The linearized equilibrium (42) can be rewritten as an eigenvalue problem:
@@
$\delta\theta^{\prime\prime}+\left(P+\mu\right)\delta\theta=0,\qquad\delta\theta(0)=0,\qquad\delta\theta^{\prime}(1)-P\epsilon\delta\theta(1)=0.$ (46)
Typically, at a supercritical bifurcation [4], the number of negative eigenvalues increases by one with each successive bifurcation. In our setup, the index of the straight equilibria (trivial solutions) for load $P_{cr,k}^{+}<P<P_{cr,k+1}^{+}$ for positive $\epsilon$ (and $P_{cr,k}^{-}<P<P_{cr,k+1}^{-}$ for negative $\epsilon$) is $k$. For a very high $\epsilon$ value, i.e., when the $\sqrt{P}\epsilon$ line approaches the vertical line, the value of $P_{cr}^{+}$ approaches zero, suggesting that bifurcations occur at a very small load when the arm is long. On the other hand, $P_{cr}^{-}$ approaches $\pi^{2}$ for higher loads, which indicates the critical loads for the case of fixed-fixed ends. The trivial solutions are stable only for loads below $P_{cr,1}^{+}$ for positive $\epsilon$ (and below $P_{cr,1}^{-}$ for negative $\epsilon$). Consider two instances of loads which correspond to first mode and second mode, say $\pi^{2}/4$ and $9\pi^{2}/4$. These chosen straight equilibria correspond to index one and two respectively, while the negative $\epsilon$ leads to index zero and index one respectively. We use these equilibria as initial solutions for numerical continuation in the subsequent sections.
@@ 4 Numerical Examples
The distinguished bifurcation diagrams facilitate the determination of the index change at a fold in the varying parameter. If there are no folds, the stability remains the same for all solutions along the branch. At folds, the stability transitions usually occur, and they were evaluated by examining the manner in which the fold occurs, as depicted in Figure 1. We use the straight trivial solutions from the previous section as initial solutions to generate the family of equilibria by varying parameters. Subsequently, we assess the stability of the resulting equilibria. Numerical continuation is performed using AUTO-07p [39], which employs pseudo-arclength continuation [1], and is capable of detecting singularities like folds and bifurcations as the parameter is varied. In these examples, the analysis is restricted solely to simple folds.